OpenStudy (anonymous):

@kittyxie1 If f(x) = 3x – 5, then what is f–1(x) ? can you help

OpenStudy (jdoe0001):

if f(x) = 3x – 5, or y = 3x – 5 swap the letters and what do you get?

OpenStudy (anonymous):

explain f-1(x)?

OpenStudy (anonymous):

its f^-1

OpenStudy (jdoe0001):

\(\huge f^{-1} \implies\) inverse function

OpenStudy (anonymous):

so what do i flip?

OpenStudy (anonymous):

inverse function y=(x+5)/3

OpenStudy (jdoe0001):

so, if you swap the variables, what would you get?

OpenStudy (jdoe0001):

no flip, just swap them out

OpenStudy (anonymous):

exchange x and y so it is x=3y-5, plus 5 on both sides,

OpenStudy (jhannybean):

First set the equation = y y = 3x - 5 replace x and y in this function x= 3y - 5 Now solve for y again. x + 5 = 3y y = (x + 5)/ 3 Rewrite it as inverse. f^{-1}(x) = (x + 5)/3

OpenStudy (anonymous):

ohhhh . thank you alll ! . Now im really bad at graphing and i dont know hot to. and the qestion is graph f(x) = –x4 + 2.

OpenStudy (jhannybean):

f(x) = -x^(4) + 2?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

f(x) = 3x – 5 y=3x-5 f–1(x) x=3y-5

OpenStudy (anonymous):

now find y

OpenStudy (jhannybean):

http://www.sketchtoy.com/ @Lesha . First graph y=x^4 + 2 by finding the x and y intercepts and then flip the graph over.

OpenStudy (anonymous):

i dont know how to .

OpenStudy (anonymous):

x=y^4+2

OpenStudy (jhannybean):

@angie041397 are you trying to show her how to find the inverse of the graph?

OpenStudy (anonymous):

yeah thats how u do the inerse in a equation